OpenStudy (anonymous):

In a direct variation, y is negative six when x is twenty-four. What is y when x equals sixteen? negative sixty-four negative fourteen negative four negative two

OpenStudy (anonymous):

@walac you know?

OpenStudy (anonymous):

Thinking about it...

OpenStudy (anonymous):

@walac K

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

@walac thank you!

OpenStudy (anonymous):

-64. you do the exact same process as what i said before.

OpenStudy (anonymous):

wait... @walac how did you get that?

OpenStudy (anonymous):

@tylerj37 how did you get that?

OpenStudy (anonymous):

divide 24 by 6. then multiply the answer by 16.

OpenStudy (anonymous):

24/-6 = 16/x, just solved the equation...

OpenStudy (anonymous):

-4*16=-64

OpenStudy (anonymous):

@tylerj37 & @walac Im confused with you too but I think im going with -4 (walac)...

OpenStudy (theeric):

Direct variation means\[\frac{y_1}{x_1}=\frac{y_2}{x_2}\] \[y_1=-6\]\[x_1=24\] \[x_2=16\] \[\frac{y_1}{x_1}=\frac{y_2}{x_2}\Rightarrow y_2=\frac{y_1\ x_2}{x_1}\]